OpenStudy (anonymous):

Can anyone tell me how to do this? w-4(4-w)=-2(2w-1)

OpenStudy (anonymous):

multiply out all brackets, collect w's on one side, numbers on the other, divide the factor multiplying w

OpenStudy (anonymous):

I did that but I keep getting the wrong answer w-16+4w=4w-2 5w-4w -16 =2 9w-16+16=9w 2=16+18 divide by 9 =2??

OpenStudy (anonymous):

first line, -4w on the righ hand side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you I think I have it now

OpenStudy (anonymous):

:)